A firm in a purely competitive industry is currently producing 1,000 units per day at a total cost of $450. If the firm produced 800 units per day, its total cost would be $300, and if it produced 500 units per day, its total cost would be $275. What are the firm’s ATC per unit at these three levels of production?

 

    At 1,000 units per day, ATC = $  .

 

    At 800 units per day, ATC = $  .

 

    At 500 units per day, ATC = $  .

 


1
Expert's answer
2020-10-16T09:32:16-0400

ATC=TC/Q

At 1,000 units per day, ATC =450/1000=$ 0.45

At 800 units per day, ATC=300/800=$0.375

At 500 units per day, ATC = 275/500=$0.55